# Adiabatic modulation of driving protocols in periodically driven quantum systems

### v1

@@ IV.3 Time scales
The problem at hand presents us with three time scales. Firstly, we have the time scale $\omega_{0}^{-1}$ set by the energy spacing of the unperturbed Hamiltonian ^$\widehat{\mathcal{H}}_{0}$. Then, we have the time period of the high frequency driving $\omega^{-1}$ and finally we have the adiabatic time scale $\Omega^{-1}$. In order for these timescales to be decoupled from each other, we would expect that $\omega\gg\omega_{o}\gg\Omega$. For a quantitative estimate of these timescales in the problem under consideration, we resort to the various energy scales in the problem.
We have already seen that $\omega_{0}$ is set by the matrix elements $\langle m|\widehat{\mathcal{H}}_{0}|n\rangle$, where $|m\rangle,|n\rangle$ are eigenstates of ^$\widehat{\mathcal{H}}_{0}$. The fast time scale is set by the average energy pumped in one fast driving period
@@
where, $|\psi\rangle$ is some general state of the system. The slow adiabatic time scale, likewise, is set by the time it takes for the parameters to change in the parameter space along some specified path $\vec{R}(\tau)=(\alpha(\tau),\beta(\tau))$ and is given by
@@
If we consider the protocol in Eq. (4), with parameters $(c_{1},c_{2},c_{3},c_{4})=(1,1,1,1)$ and start the time evolution from an initial point $(\alpha\left(\tau_{0}\right),\beta\left(\tau_{0}\right))$ in the parameter space, we have $\Delta E_{\text{fast}}=\langle\psi|\mathbf{J}\cdot\mathcal{B}_{f}|\psi\rangle$, where
@@
It follows from the conservation of energy that the energy cost for the slow evolution $\Delta E_{\text{slow}}$ is a large multiple of the energy cost per fast driving cycle $\Delta E_{\text{fast}}$. Thus, the adiabatic condition on the frequencies $\Omega$ and $\omega$ can be obtained from the inequality, $\Delta E_{\text{fast}}\ll\Delta E_{\text{slow}}$.
@@ IV.4 Summary
In this work, we have formulated a method to adiabatically modulate driving protocols in periodically driven quantum systems. We began with a parametrized fast-driving protocol and computed the perturbative corrections to the effective Hamiltonian and the micro-motion kick operators. Then, we induce a second, much slower drive by an adiabatic variation of these parameters. We study the effects of this variation through a simple rigid rotor system. We note the presence of various diabolical points and diabolical loci in the spectrum of this Hamiltonian. We further deduce various properties of the eigenspace of this Hamiltonian through its connection to the parameter space by constructing various adiabatic paths. We observe as a consequence of the above connection that a small deformation to the path in the parameter space could yield a new path in the Hilbert space that is not homeomorphic to the original path on the Hilbert space. Such small deformations that have negligible energy cost may thereby bring about discrete changes in the topological properties of the eigenstates. We also provide a generalization of our parametrization of the driving protocol, which could potentially yield a much more enriched energy spectrum and distribution of degeneracies in the parameter space. Finally, we comment on the regime where our analysis is valid by comparing the different energy scales involved.
We conclude by noting that the method developed in this work can be applied to any periodically driven system, whereby the topological aspects pointed out by us can find diverse ramifications in theoretical or experimental contexts.
@@ V Acknowledgments
The authors would like to thank Aritra Banerjee and Pritam K Jana for useful discussions and suggestions. A.M. would like to thank Abhirath Anand for help with Julia. Figures were created using Makie.jl and Inkscape.

### v2

@@ IV.3 Time scales
The problem at hand presents us with three time scales. Firstly, we have the time scale $\omega_{0}^{-1}$ set by the energy scale of the unperturbed Hamiltonian ^$\widehat{\mathcal{H}}_{0}$. Then, we have the time period of the high frequency driving $\omega^{-1}$ and finally we have the adiabatic time scale $\Omega^{-1}$. In order for these timescales to be decoupled from each other, we would expect that $\omega\gg\omega_{o}\gg\Omega$. For a quantitative estimate of these timescales in the problem under consideration, we resort to the various energy scales in the problem.
We have already seen that $\omega_{0}$ is set by the matrix elements $\langle m|\widehat{\mathcal{H}}_{0}|n\rangle$, where $|m\rangle,|n\rangle$ are eigenstates of ^$\widehat{\mathcal{H}}_{0}$. Our formalism rests upon the existence of a suitable timescale (high frequency driving) such that the original time dependent problem reduces to an effective time independent problem. The frequency $\omega$ has to be chosen with a tolerance set by the validity of Eq. (6). Averaging over such time scales $\left(\sim\frac{1}{\omega}\right)$ thus gives the effective averaged contribution of the potential within a period.
@@
where, $|\psi\rangle$ is some general state of the system. The effective Hamiltonian takes into account the averaging over the fast time period. If we allow the ^$\widehat{\mathcal{H}}_{\text{eff}}$ to undergo slow modulation owing to slow changes in $\alpha,\beta$ then under the assumption that the two timescales are widely separated and therefore decoupled. we can, define a slow energy scale $\Delta E_{\text{slow}}$. The slow adiabatic time scale, is set by the time it takes for the parameters to change in the parameter space along some specified path $\vec{R}(\tau)=(\alpha(\tau),\beta(\tau))$ and is given by the integral of $\Braket{\frac{d\widehat{\mathcal{H}}_{\text{eff}}\left(R(\tau)\right)}{d\tau}}$ over the adiabatic timescale. Thus
@@
If we consider the protocol in Eq. (4), with parameters $(c_{1},c_{2},c_{3},c_{4})=(1,1,1,1)$ and start the time evolution from an initial point $(\alpha\left(\tau_{0}\right),\beta\left(\tau_{0}\right))$ in the parameter space, we have $\Delta E_{\text{fast}}=\langle\psi|\mathbf{S}\cdot\mathcal{B}_{f}|\psi\rangle$, where
@@
It follows from the conservation of energy that the energy cost for the slow evolution $\Delta E_{\text{slow}}$ is a large multiple of the typical energy cost per fast driving cycle $\Delta E_{\text{fast}}$. Thus, the adiabatic condition on the frequencies $\Omega$ and $\omega$ can be obtained from the inequality, $\Delta E_{\text{fast}}\ll\Delta E_{\text{slow}}$.
@@ IV.4 Summary
In this work, we have formulated a method to adiabatically modulate driving protocols in periodically driven quantum systems. We began with a parametrized fast-driving protocol and computed the perturbative corrections to the effective Hamiltonian and the micro-motion kick operators. Then, we induce a second, much slower drive by an adiabatic variation of these parameters. We study the effects of this variation through a simple spin $1/2$ system. We note the presence of various diabolical points and diabolical loci in the spectrum of this Hamiltonian. We further deduce various properties of the eigenspace of this Hamiltonian through its connection to the parameter space by constructing various adiabatic paths. We emphasise that while adiabatic evolution in Hamiltonians can be brought about in numerous ways, our formulation of using the temporal partitions of the driving protocol as the parameters allows for arbitrarily large and complex parameter spaces without requiring additional couplings or operators. It is also important to note that the structure of degeneracy is related to the zeros of certain polynomials of these partitioning parameters and not the properties of the operators involved in the driving protocol. This has the potential to adjust energy gaps in physically relevant systems simply through tuning the parameters and/or increasing the number of parameters. While we have worked with only two parameters, we have provided a generalization of our parametrization of the driving protocol, which could potentially yield a much more enriched energy spectrum and distribution of degeneracies in the parameter space. We observe that as a consequence of such degeneracy structures, a small deformation to the path in the parameter space could yield a new path in the Hilbert space that is not homeomorphic to the original path on the Hilbert space. Such small deformations have a negligible energy cost for the driving agency. However, they bring about discrete changes in the topological properties (crossings) of the eigenstates. In the simplest case of the spin-1/2 system, such crossings imply that for small changes in the cyclic path in the parameter space, the direction of the synthesized magnetic field may loop back to itself multiple times, which may be physically significant. We also note that if the potentials themselves were related to, say, a lattice model or if the parameters $\alpha,\beta$ could be interpreted in the Fourier space as $k_{x},k_{y}$ whereby the energy spectrum becomes a band structure in the Brillouin zone, then the rich structure of the diabolical points and loci in that context may actually lead to physically observable properties by choosing appropriate trajectories on the parameter space. Finally, we comment on the regime where our analysis is valid by comparing the different energy scales involved. We conclude by noting that the method developed in this work can be applied to any periodically driven system, whereby the topological aspects pointed out by us can find diverse ramifications in theoretical or experimental contexts.
@@ V Acknowledgments
The authors would like to thank Aritra Banerjee and Pritam K. Jana for useful discussions and suggestions. A.M. would like to thank Abhirath Anand for help with Julia. Figures were created using Makie.jl and Inkscape.